A solar collector is 2.5m long by 2.0m wide. It is held in place by a frame of uniform width around its outside edge. If the exposed collector area is 2.5m^2 , what is the width of the frame, to the nearest tenth of a centimeter?
Answer by stanbon(75887)

A solar collector is 2.5m long by 2.0m wide. It is held in place by a frame of uniform width around its outside edge. If the exposed collector area is 2.5m^2 , what is the width of the frame, to the nearest tenth of a centimeter
----------------------------
Draw the picture.
If the width of the frame is x the dimensions of the exposed panel are
length = 2.5-2x and width is 2-2x
EQUATION:
length * width = 2.5 m^2
(2.5-2x)(2-2x)=2.5
(2.5-2x)(1-x)= 1.25
2.5-2.5x-2x+2x^2=1.25
2x^2-4.5x+1.25=0
x=[4.5+-sqrt(4.5^2-4(2)(1.25)]/4
x=[4.5+-sqrt3.20156212..]/4
x=1.5723... or x=0.67767726...
Only x=0.67767726 is a viable answer as the other value
would give a 2x value greater than the width of length of the panel.
Cheers,
Stan H.